David is going to purchase two stocks to form the initial holdings in his portfolio. Iron stock has an expected return of 15 percent, while Copper stock has an expected return of 20 percent.

If David plans to invest 30 percent of his funds in Iron and the remainder in Copper, what will be the expected return from his portfolio? (Round answer to 1 decimal place, e.g. 17.5%.)
Expected return

%
What would the expected return of David's portfolio invests 70 percent of his funds in Iron stock? (Round answer to 1 decimal place, e.g. 17.5%.)

Expected return

%
Answer #1

Solution :-

Expected Return from Iron Stock = 15%

Expected Return from Copper Stock = 20%

now

(A)

If Weights of funds invested in Iron = 30% = 0.30

Then Weights of funds invested in Copper = ( 100 - 30 ) = 70% = 0.70

Now Expected Return of Portfolio = ( Weight of Iron * ER of Iron ) + ( Weight of Copper * ER of Copper )

= ( 0.30 * 15% ) + ( 0.70 * 20 )

= 4.5% + 14%  

= 18.5%

(B)

If Weights of funds invested in Iron = 70% = 0.70

Then Weights of funds invested in Copper = ( 100 - 70 ) = 30% = 0.30

Now Expected Return of Portfolio = ( Weight of Iron * ER of Iron ) + ( Weight of Copper * ER of Copper )

= ( 0.70 * 15% ) + ( 0.30 * 20 )

= 10.5% + 6%  

= 16.5%

If there is any doubt please ask in comments